If the market demand decreases by 50% what will be the new clearing price and


quantity for this market?

1
Expert's answer
2022-02-20T17:02:59-0500

A decrease in demand will cause the equilibrium price to fall; the quantity supplied will decrease.

The clearing price is also the equilibrium price. This is where the quantity of demand in supply is equal to the number of supplies available. We can use the linear algebraic expression to find the supply line and demand line of a product.

linear supply line is;

The linear supply function is:

Qs = x + yP

Where:

  • Qs = the quantity supplied
  • X = quantity
  • P = price

The linear demand function is:

Qd = x + yP

Where:

  • Qd = the quantity of demand
  • X = quantity
  • P = price

The equilibrium price sets the two equal to each other:

Qs = Qd.

if demand decreases by 50% then the new clearing price will be;

"\frac {x*50}{100}" =0.5x and price will be 0.5p.

Qs = Qd=x-0.5x=p-0.5p